# Actions Speak Louder than Words: Trillion-Parameter Sequential Transducers for Generative Recommendations

### v1

@@ Abstract
Machine Learning, Recommendations, Generative Modeling, Generative Recommenders
@@ 1 Introduction
Recommendation systems, quintessential in the realm of online content platforms and e-commerce, play a pivotal role in personalizing billions of user experiences on a daily basis. State-of-the-art approaches in recommendations have been based on Deep Learning Recommendation Models (DLRMs) (Mudigere et al., 2022) for about a decade (Covington et al., 2016; Cheng et al., 2016; Zhou et al., 2018; Tang et al., 2020; Wang et al., 2021; Xia et al., 2023). DLRMs are characterized by their usage of heterogeneous features, such as numerical features – counters and ratios, embeddings, and categorical features such as creator ids, user ids, etc. Due to new content and products being added every minute, the feature space is of extreme high cardinality, often in the range of billions (Eksombatchai et al., 2018). To leverage tens of thousands of such features, DLRMs employ various neural networks to combine features, transform intermediate representations, and finally compose the final outputs.
@@
Inspired by the success achieved by Transformers in language and vision, we revisit fundamental design choices in modern recommendation systems. We observe that alternative formulations at billion-user scale need to overcome three challenges. First, features in recommendation systems lack explicit structures. While sequential formulations have been explored in small-scale settings (Kang & McAuley, 2018; Sun et al., 2019; Zhou et al., 2020), heterogeneous features, including high cardinality ids, cross features, counters, ratios, etc. play critical roles in industry-scale DLRMs (Mudigere et al., 2022). Second, recommendation systems use billion-scale vocabularies that change continuously. A billion-scale dynamic vocabulary, in contrast to 100K-scale static ones in language (Brown et al., 2020), creates training challenges and necessitates high inference cost given the need to consider tens of thousands of candidates in a target-aware fashion (Zhou et al., 2018; Wang et al., 2020). Finally, computational cost represents the main bottleneck in enabling large-scale sequential models. GPT-3 was trained on a total of 300B tokens over a period of 1-2 months with thousands of GPUs (Brown et al., 2020). This scale appears daunting, until we contrast it with the scale of user actions. The largest internet platforms serve billions of daily active users, who engage with billions of posts, images, and videos per day. User sequences could be of length up to $10^{5}$ in extreme cases (Chang et al., 2023). Consequentially, recommendation systems need to handle a few orders of magnitude more tokens per day than what language models process over 1-2 months.
@@
We next address computational cost challenges throughout training and inference. We propose a new sequential transduction architecture, Hierarchical Sequential Transduction Units (HSTU). HSTU modifies attention mechanism for large, non-stationary vocabulary, and exploits characteristics of recommendation datasets to achieve 5.3x to 15.2x speedup vs FlashAttention2-based Transformers on 8192 length sequences. Further, through a new algorithm, M-FALCON, that fully amortizes computational costs via micro-batching (Section 3.4), we can serve 285x more complex GR models while achieving 1.50x-2.48x speedups, all with the same inference budget used by traditional DLRMs.
@@ 2.1 Unifying heterogeneous feature spaces in DLRMs
Figure 2: Comparison of features and training procedures: DLRMs vs GRs. $E,F,G,H$ denote categorical features. $\Phi_{i}$ represents the $i$-th item in the merged main time series. $\Psi_{i}(t_{j})$ denotes training example $i$ emitted at time $j$.
@@ 2.2 Reformulating ranking and retrieval as sequential transduction tasks
Given a list of tokens $x_{0},x_{1},\ldots,x_{n-1}$ ordered chronologically, the time when those tokens are observed $t_{0},t_{1},\ldots,t_{n-1}$, and other metadata (such as user actions on those tokens $a_{i}$s, if $x_{i}$ is a token from the main time series discussed in Section 2.1), a sequential transduction task maps those input sequences to the output tokens $y_{0},y_{1},\ldots,y_{n-1}$ subject to a mask sequence $m_{0},m_{1},\ldots,m_{n-1}$ ($m_{i}\in\{0,1\}$), where $m_{i}=0$ indicates that $y_{i}$ is undefined. The input tokens come from a dynamic, non-stationary vocabulary $\mathbb{X}$ where the contents (e.g., videos that users engage with) are $\mathbb{X}_{c}\subseteq\mathbb{X}$. Full notations can be found in Appendix A. We consider causal masked autoregressive settings for these tasks throughout the rest of this section.
Retrieval. In GRs, retrieval tasks learn a distribution $p(x_{i+1}|u_{i})$ for each user over $x_{i+1}\in\mathbb{X}_{c}$, where $u_{i}$ is the user’s representation at step $i$. A typical objective is to select $\operatorname*{arg\,max}_{x\in\mathbb{X}_{c}}p(x|u_{i})$ to maximize a specific reward. This differs from a standard autoregressive setup in two ways. First, the supervision for $x_{i}$, $y_{i}$, is not necessarily $x_{i+1}$, as users could respond negatively to $x_{i+1}$. Second, $y_{i}$ is undefined in situations where the next token represents a non-engagement related categorical feature, such as demographics ($x_{i+1}\notin\mathbb{X}_{c}$). For these cases, we set $m_{i}=0$.
Ranking. Ranking tasks in GRs pose unique challenges as modern recommendation systems often require a “target-aware” formulation. In such a formulation, “interaction” of target, $x_{i+1}$, and historical features up to $i$ needs to occur as early as possible, which is infeasible with a standard autoregressive setup where “interaction” typically happens late (e.g., via softmax after encoder output). We address this problem by interleaving items and actions in the main time series. The resulting new time series (before categorical features $x\notin\mathbb{X}_{c}$) is then $x_{0},a_{0},x_{1},a_{1},\ldots,x_{n-1},a_{n-1}$, where mask $m_{i}$s are 00s for the action positions. We apply a small neural network to transform predictions at content positions into multi-task predictions in practice. This approach enables us to apply target-aware cross-attention to all $n$ (user, item) engagements in one pass with causal masking.
@@ 3 A High Performance Self-Attention Encoder for Generative Recommendations
In order to scale up GRs for industrial-scale recommendation systems with extremely large, non-stationary vocabularies and massive amount of data, we next introduce a new encoder design for GRs called Hierarchical Sequential Transduction Unit (HSTU). HSTU is composed of a stack of identical layers connected by residual connections (He et al., 2015). Each layer consists of three main sub-layers: Pointwise Projection (Equation 1), Spatial Aggregation (Equation 2), and Pointwise Transformation (Equation 3):
@@
where $f_{i}(X)$ denotes an MLP; we use one linear layer, $f_{i}(X)=W_{i}(X)+b_{i}$ for $f_{1}$ and $f_{2}$ to reduce compute complexity; $\phi_{1}$ and $\phi_{2}$ denote nonlinearity, for both of which we use SiLU (Elfwing et al., 2017); and $\text{rab}^{p,t}$ denotes relative attention bias (Raffel et al., 2020) that incorporates positional ($p$) and temporal ($t$) information.
@@
Feature Interaction is the most critical part of DLRMs. Common approaches used include factorization machines and their neural network variants  (Rendle, 2010; Guo et al., 2017; Xiao et al., 2017), higher order feature interactions (Wang et al., 2021), and their deep residual variants (Zhang et al., 2022). HSTU replaces feature interactions by enabling attention pooled features to directly “interact” with other features via $\text{Norm}\left(A(X)V(X)\right)\odot U(X)$. This design is motivated by the difficulty of approximating dot products with learned MLPs (Rendle et al., 2020; Zhai et al., 2023a). Given SiLU is applied to $U(X)$, $\text{Norm}\left(A(X)V(X)\right)\odot U(X)$ can also be interpreted as a variant of SwiGLU (Shazeer, 2020).
@@ 3.1 Pointwise aggregated attention
HSTU adopts a new pointwise aggregated attention mechanism instead of softmax attention in Transformers. This is motivated by two factors. First, in recommendations, the number of prior data points related to target serves as a strong feature indicating the intensity of user preferences, which is hard to capture after the softmax normalization. This is critical as we need to predict both the intensity of engagements, e.g., time spent on a given item, and the relative ordering of the items, e.g., predicting an ordering of candidates to maximize AUC. Second, while softmax activation is robust to noise by construction, it is less suited for non-stationary vocabularies in streaming settings.
The proposed pointwise aggregated attention mechanism is depicted in Equation 2. Importantly, layer norm is needed after pointwise pooling to stabilize training. One way to understand this design is through synthetic data following a Dirichlet Process that generates streaming data over a nonstationary vocabulary (details in Appendix B). In this setting, we can observe gaps as large as 44.7% between softmax and pointwise attention setups as shown in Table 1.
@@ 3.2 Leveraging and algorithmically increasing sparsity
In recommendation systems, the length of user history sequences often follows a skewed distribution, leading to sparse input sequences, particularly in the settings with very long sequences. This sparsity can be leveraged to significantly improve the efficiency of the encoder. To achieve this, we have developed an efficient attention kernel for GPUs that fuses back-to-back GEMMs in a manner similar to FlashAttention (Rabe & Staats, 2021; Dao et al., 2022) but performs fully raggified attention computations. This essentially transforms the attention computation into grouped GEMMs of various sizes. As a result, self-attention in HSTU becomes memory-bound and scales as $\Theta(\sum_{i}n_{i}^{2}d_{qk}^{2}R^{-1})$ in terms of memory accesses, where $n_{i}$ is the sequence length for sample $i$, $d_{qk}$ is attention dimension, and $R$ is the register size. This approach by itself leads to 2-5x throughput gains as discussed in Section 4.2.
@@
Specifically, let $\Gamma(n,L)$ be a function that selects a subsequence of length $L$ from the original sequence $x_{0},\ldots,x_{n-1}$. $SL$ selects input sequences as follows:
$\small\begin{split}x_{0},\ldots,x_{n_{i}-1}&\text{ if }n_{i}\leq N^{\alpha/2}\\ \Gamma(n_{i},N^{\alpha/2})&\text{ if }n_{i}>N^{\alpha/2},\text{w/ probability }1-N^{\alpha}/n_{i}^{2}\\ x_{0},\ldots,x_{n_{i}-1}&\text{ if }n_{i}>N^{\alpha/2},\text{w/ probability }N^{\alpha}/n_{i}^{2}\end{split}$ (4)
which reduces overall attention-related complexity to $O(N^{\alpha}d)$ for $\alpha\in(1,2]$. We remark that applying SL to training leads to an overall cost-effective model system design, as training stage involves a significantly higher computational cost compared to the inference stage.
Table 2 presents the sparsity (see Appendix D) for different sequence lengths and $\alpha$ values, for a representative industry-scale configuration with 30-day user history. The settings that result in negligible regression in model quality are underlined and highlighted in blue. The rows labeled “$\alpha=2.0$” represents the base sparsity case where SL is not applied. Lower $\alpha$’s are applicable to longer sequences up to the longest sequence length we tested, 8,192.
@@ 3.4 Scaling up inference via cost-amortization
Within a forward pass, M-FALCON handles $b_{m}$ candidates in parallel by modifying attention masks and $\text{rab}^{p,t}$ biases such that the attention operations performed for $b_{m}$ candidates are exactly the same. This reduces the cost of applying cross-attention from $O(b_{m}n^{2}d)$ to $O((n+b_{m})^{2}d)=O(n^{2}d)$ when $b_{m}$ can be considered a small constant relative to $n$. We optionally divide the overall $m$ candidates into $\lceil m/b_{m}\rceil$ microbatches of size $b_{m}$ to leverage encoder-level KV caching (Pope et al., 2022) either across forward passes to reduce cost, or across requests to minimize tail latency.
Overall, M-FALCON enables model complexity to linearly scale up with the number of candidates in traditional DLRMs’s ranking stages; we succeeded in applying a 285x more complex target-aware cross attention model at 1.5x throughput with a constant inference budget for a typical ranking configuration discussed in Section 4.3.

### v2

@@ Abstract
Generative Recommenders, Machine Learning, Recommendations, Generative Modeling, Sequential Recommendations, HSTU, ICML
@@ 1 Introduction
Recommendation systems, quintessential in the realm of online content platforms and e-commerce, play a pivotal role in personalizing billions of user experiences on a daily basis. State-of-the-art approaches in recommendations have been based on Deep Learning Recommendation Models (DLRMs) (Mudigere et al., 2022) for about a decade (Covington et al., 2016; Cheng et al., 2016; Zhou et al., 2018; Tang et al., 2020; Wang et al., 2021; Xia et al., 2023). DLRMs are characterized by their usage of heterogeneous features, such as numerical features – counters and ratios, embeddings, and categorical features such as creator ids, user ids, etc. Due to new content and products being added every minute, the feature space is of extreme high cardinality, often in the range of billions (Eksombatchai et al., 2018). To leverage tens of thousands of such features, DLRMs employ various neural networks to combine features, transform intermediate representations, and compose the final outputs.
@@
Inspired by the success achieved by Transformers in language and vision, we revisit fundamental design choices in modern recommendation systems. We observe that alternative formulations at billion-user scale need to overcome three challenges. First, features in recommendation systems lack explicit structures. While sequential formulations have been explored in small-scale settings (detailed discussions in Appendix B), heterogeneous features, including high cardinality ids, cross features, counters, ratios, etc. play critical roles in industry-scale DLRMs (Mudigere et al., 2022). Second, recommendation systems use billion-scale vocabularies that change continuously. A billion-scale dynamic vocabulary, in contrast to 100K-scale static ones in language (Brown et al., 2020), creates training challenges and necessitates high inference cost given the need to consider tens of thousands of candidates in a target-aware fashion (Zhou et al., 2018; Wang et al., 2020). Finally, computational cost represents the main bottleneck in enabling large-scale sequential models. GPT-3 was trained on a total of 300B tokens over a period of 1-2 months with thousands of GPUs (Brown et al., 2020). This scale appears daunting, until we contrast it with the scale of user actions. The largest internet platforms serve billions of daily active users, who engage with billions of posts, images, and videos per day. User sequences could be of length up to $10^{5}$ (Chang et al., 2023). Consequentially, recommendation systems need to handle a few orders of magnitude more tokens per day than what language models process over 1-2 months.
@@
We next address computational cost challenges throughout training and inference. We propose a new sequential transduction architecture, Hierarchical Sequential Transduction Units (HSTU). HSTU modifies attention mechanism for large, non-stationary vocabulary, and exploits characteristics of recommendation datasets to achieve 5.3x to 15.2x speedup vs FlashAttention2-based Transformers on 8192 length sequences. Further, through a new algorithm, M-FALCON, that fully amortizes computational costs via micro-batching (Section 3.4), we can serve 285x more complex GR models while achieving 1.50x-2.99x speedups, all with the same inference budget used by traditional DLRMs.
@@ 2.1 Unifying heterogeneous feature spaces in DLRMs
Figure 2: Comparison of features and training procedures: DLRMs vs GRs. $E,F,G,H$ denote categorical features. $\Phi_{i}$ represents the $i$-th item in the merged main time series. $\Psi_{k}(t_{j})$ denotes training example $k$ emitted at time $t_{j}$. Full notations can be found in Appendix A.
@@ 2.2 Reformulating ranking and retrieval as sequential transduction tasks
Given a list of $n$ tokens $x_{0},x_{1},\ldots,x_{n-1}$ ($x_{i}\in\mathbb{X}$) ordered chronologically, the time when those tokens are observed $t_{0},t_{1},\ldots,t_{n-1}$, a sequential transduction task maps this input sequence to the output tokens $y_{0},y_{1},\ldots,y_{n-1}$ ($y_{i}\in\mathbb{X}\cup\{\varnothing\}$), where $y_{i}=\varnothing$ indicates that $y_{i}$ is undefined.
We use $\Phi_{i}\in\mathbb{X}_{c}$ ($\mathbb{X}_{c}\subseteq\mathbb{X}$) to denote a content (e.g., images or videos) that the system provides to the user. Given new content are constantly created, $\mathbb{X}_{c}$ and $\mathbb{X}$ are non-stationary. The user can respond to $\Phi_{i}$ with some action $a_{i}$ (e.g., like, skip, video completion+share) $a_{i}\in\mathbb{X}$. We denote the total number of contents that a user has interacted with by $n_{c}$.
The standard ranking and retrieval tasks, in causal autoregressive settings, can then be defined as sequential transduction tasks (Table 1). We make the following observations:
Task Specification (Inputs / Outputs)
Ranking $x_{i}$s $\Phi_{0},a_{0},\Phi_{1},a_{1},\ldots,\Phi_{n_{c}-1},a_{n_{c}-1}$
$y_{i}$s $a_{0},\varnothing,a_{1},\varnothing,\ldots,a_{n_{c}-1},\varnothing$
Retrieval $x_{i}$s $(\Phi_{0},a_{0}),(\Phi_{1},a_{1}),\ldots,(\Phi_{n_{c}-1},a_{n_{c}-1})$
$y_{i}$s $\Phi_{1}^{\prime},\Phi_{2}^{\prime},\ldots,\Phi_{n_{c}-1}^{\prime},\varnothing$
($\Phi_{i}^{\prime}=\Phi_{i}$ if $a_{i}$ is positive, otherwise $\varnothing$)
Table 1: Ranking and retrieval as sequential transduction tasks. Other categorical features are omitted for simplicity. We compare GRs with traditional sequential recommenders in Section B.2.
Retrieval. In recommendation system’s retrieval stage, we learn a distribution $p(\Phi_{i+1}|u_{i})$ over $\Phi_{i+1}\in\mathbb{X}_{c}$, where $u_{i}$ is the user’s representation at token $i$. A typical objective is to select $\operatorname*{arg\,max}_{\Phi\in\mathbb{X}_{c}}p(\Phi|u_{i})$ to maximize some reward. This differs from a standard autoregressive setup in two ways. First, the supervision for $x_{i}$, $y_{i}$, is not necessarily $\Phi_{i+1}$, as users could respond negatively to $\Phi_{i+1}$. Second, $y_{i}$ is undefined when $x_{i+1}$ represents a non-engagement related categorical feature, such as demographics.
Ranking. Ranking tasks in GRs pose unique challenges as industrial recommendation systems often require a “target-aware” formulation. In such settings, “interaction” of target, $\Phi_{i+1}$, and historical features needs to occur as early as possible, which is infeasible with a standard autoregressive setup where “interaction” happens late (e.g., via softmax after encoder output). We address this by interleaving items and actions in Table 1, which enables the ranking task to be formulated as $p(a_{i+1}|\Phi_{0},a_{0},\Phi_{1},a_{1},\ldots,\Phi_{i+1})$ (before categorical features). We apply a small neural network to transform outputs at $\Phi_{i+1}$ into multi-task predictions in practice. Importantly, this enables us to apply target-aware cross-attention to all $n_{c}$ engagements in one pass.
@@ 3 A High Performance Self-Attention Encoder for Generative Recommendations
To scale up GRs for industrial-scale recommendation systems with large, non-stationary vocabularies, we next introduce a new encoder design, Hierarchical Sequential Transduction Unit (HSTU). HSTU consists of a stack of identical layers connected by residual connections (He et al., 2015). Each layer contains three sub-layers: Pointwise Projection (Equation 1), Spatial Aggregation (Equation 2), and Pointwise Transformation (Equation 3):
@@
where $f_{i}(X)$ denotes an MLP; we use one linear layer, $f_{i}(X)=W_{i}(X)+b_{i}$ for $f_{1}$ and $f_{2}$ to reduce compute complexity and further batches computations for queries $Q(X)$, keys $K(X)$, values $V(X)$, and gating weights $U(X)$ with a fused kernel; $\phi_{1}$ and $\phi_{2}$ denote nonlinearity, for both of which we use SiLU (Elfwing et al., 2017); Norm is layer norm; and $\text{rab}^{p,t}$ denotes relative attention bias (Raffel et al., 2020) that incorporates positional ($p$) and temporal ($t$) information. Full notations used can be found in Table 9.
@@
Feature Interaction is the most critical part of DLRMs. Common approaches used include factorization machines and their neural network variants  (Rendle, 2010; Guo et al., 2017; Xiao et al., 2017), higher order feature interactions (Wang et al., 2021), etc. HSTU replaces feature interactions by enabling attention pooled features to directly “interact” with other features via $\text{Norm}\left(A(X)V(X)\right)\odot U(X)$. This design is motivated by the difficulty of approximating dot products with learned MLPs (Rendle et al., 2020; Zhai et al., 2023a). Given SiLU is applied to $U(X)$, $\text{Norm}\left(A(X)V(X)\right)\odot U(X)$ can also be interpreted as a variant of SwiGLU (Shazeer, 2020).
@@ 3.1 Pointwise aggregated attention
HSTU adopts a new pointwise aggregated (normalized) attention mechanism (in contrast, softmax attention computes normalization factor over the entire sequence). This is motivated by two factors. First, the number of prior data points related to target serves as a strong feature indicating the intensity of user preferences, which is hard to capture after softmax normalization. This is critical as we need to predict both the intensity of engagements, e.g., time spent on a given item, and the relative ordering of the items, e.g., predicting an ordering to maximize AUC. Second, while softmax activation is robust to noise by construction, it is less suited for non-stationary vocabularies in streaming settings.
The proposed pointwise aggregated attention mechanism is depicted in Equation 2. Importantly, layer norm is needed after pointwise pooling to stabilize training. One way to understand this design is through synthetic data following a Dirichlet Process that generates streaming data over a nonstationary vocabulary (details in Appendix C). In this setting, we can observe gaps as large as 44.7% between softmax and pointwise attention setups as shown in Table 2.
@@ 3.2 Leveraging and algorithmically increasing sparsity
In recommendation systems, the length of user history sequences often follows a skewed distribution, leading to sparse input sequences, particularly in the settings with very long sequences. This sparsity can be leveraged to significantly improve the efficiency of the encoder. To achieve this, we have developed an efficient attention kernel for GPUs that fuses back-to-back GEMMs in a manner similar to (Rabe & Staats, 2021; Dao et al., 2022) but performs fully raggified attention computations. This essentially transforms the attention computation into grouped GEMMs of various sizes (Appendix G). As a result, self-attention in HSTU becomes memory-bound and scales as $\Theta(\sum_{i}n_{i}^{2}d_{qk}^{2}R^{-1})$ in terms of memory accesses, where $n_{i}$ is the sequence length for sample $i$, $d_{qk}$ is attention dimension, and $R$ is the register size. This approach by itself leads to 2-5x throughput gains as discussed in Section 4.2.
@@
We can represent user $j$’s history as a sequence $(x_{i})_{i=0}^{n_{c,j}}$, where $n_{c,j}$ is the number of contents user interacted with. Let $N_{c}=\max_{j}{n_{c,j}}$. Let $(x_{i_{k}})_{k=0}^{L}$ be a subsequence of length $L$ constructed from the original sequence $(x_{i})_{i=0}^{n_{c,j}}$. $SL$ selects input sequences as follows:
$\small\begin{split}{(x_{i})_{i=0}^{n_{c,j}}}&\text{ if }n_{c,j}\leq N_{c}^{\alpha/2}\\ {(x_{i_{k}})_{k=0}^{N_{c}^{\alpha/2}}}&\text{ if }n_{c,j}>N_{c}^{\alpha/2},\text{w/ probability }1-N_{c}^{\alpha}/n_{c,j}^{2}\\ {(x_{i})_{i=0}^{n_{c,j}}}&\text{ if }n_{c,j}>N_{c}^{\alpha/2},\text{w/ probability }N_{c}^{\alpha}/n_{c,j}^{2}\end{split}$ (4)
which reduces attention-related complexity to $O(N_{c}^{\alpha}d)=O(N^{\alpha}d)$ for $\alpha\in(1,2]$. A more thorough discussion of subsequence selection can be found in Section F.1. We remark that applying SL to training leads to a cost-effective system design, as training generally involves a significantly higher computational cost compared to inference.
Table 3 presents the sparsity (see Appendix F) for different sequence lengths and $\alpha$ values, for a representative industry-scale configuration with 30-day user history. The settings that result in negligible regression in model quality are underlined and highlighted in blue. The rows labeled “$\alpha=2.0$” represents the base sparsity case where SL is not applied. Lower $\alpha$’s are applicable to longer sequences up to the longest sequence length we tested, 8,192.
@@ 3.4 Scaling up inference via cost-amortization
Within a forward pass, M-FALCON handles $b_{m}$ candidates in parallel by modifying attention masks and $\text{rab}^{p,t}$ biases such that the attention operations performed for $b_{m}$ candidates are exactly the same. This reduces the cost of applying cross-attention from $O(b_{m}n^{2}d)$ to $O((n+b_{m})^{2}d)=O(n^{2}d)$ when $b_{m}$ can be considered a small constant relative to $n$. We optionally divide the overall $m$ candidates into $\lceil m/b_{m}\rceil$ microbatches of size $b_{m}$ to leverage encoder-level KV caching (Pope et al., 2022) either across forward passes to reduce cost, or across requests to minimize tail latency (More detailed discussions in Appendix H).
Overall, M-FALCON enables model complexity to linearly scale up with the number of candidates in traditional DLRMs’s ranking stages; we succeeded in applying a 285x more complex target-aware cross attention model at 1.5x-3x throughput with a constant inference budget for a typical ranking configuration discussed in Section 4.3.